A planet of mass M moves around the Sun along an ellipse so that its minimum distance from the Sun is equal to r and the maximum distance to R. Making use of Kepler s laws, find its period of revolution around the Sun.
1
Expert's answer
2020-10-16T11:00:13-0400

Explanations & Calculations


  • According to Kepler's 3rd law, "\small T^2 \propto a^3" : a = the semi major axis
  • Accordingly the periodic time does not depend on the eccentricity as any time delay is caught up with the increased speeds on such location of the path.
  • Therefore, an average value for the semi major axis could be written as "\small a = \frac{(r+R)}{2}"
  • From the motion of a uniform circle by the application of Newton's second law towards center it could be written for the period, "\small T^2 = \large\frac{4\pi^2}{GM_0}r^3 \small \to T^2 \propto r^3"
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small T^2 &= \small \frac{4\pi^2}{GM_0}\times \frac{(r+R)^3}{8}\\
\small T &= \small \pi \sqrt{\frac{(r+R)^3}{2GM_0}} 
\end{aligned}" : M0 is the mass of the Sun